# Probing Visual Language Priors in VLMs

### v1

@@ 5.1 Image DPO
In this section, we examine our proposed Image-DPO (Section 4.1) on both the ViLP  benchmark and general VQA benchmarks. Specifically, for self-generating question-image-answer (QIA) pairs, we use COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017) as our seed datasets. Based on these seed datasets, we instruct the VLM to leverage pretrained image generation and editing models (Liu et al., 2023d; Podell et al., 2023; Brooks et al., 2023b) to generate new images and corresponding QA pairs. After generating the QA pairs, we corrupt the images either semantically or at the pixel level, while keeping the QA pairs unchanged, to create the good and bad QIA pairs used for Image DPO training. Details of our image generation method are provided in the Appendix C and training details are in the Appendix C.2
@@
LLaVA-1.5-7B + CSR 27.50 64.33 40.84 63.67
LLaVA-1.5-7B + RLHF-V 29.50 75.00 36.33 65.33
LLaVA-1.5-7B + Text-DPO 31.34 71.67 37.83 65.67
LLaVA-1.5-7B + Image-DPO 34.17 77.67 38.67 67.33
For comparison, we introduce another baseline, Text-DPO, which follows the same question-image-answer generation process as our Image-DPO but applies LLM self-rewarding techniques as outlined in (Yuan et al., 2024). In Text-DPO, good and bad pairs are created by sampling positive and negative answers generated by the VLM, while keeping the image and question unchanged. The difference between Image-DPO and Text-DPO is illustrated in Figure 6. The green box corresponds to Image-DPO, where multiple corrupted images are generated through semantic alterations, Gaussian blurring, and pixelation, while the question and answer remain constant. In contrast, Text-DPO (purple-box) maintains the same image and generates multiple answers, each associated with a rating. It is worth noting that while Text-DPO shares similarities with other VLM self-rewarding techniques (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a), a key distinction is our use of pre-trained image generation models to create diverse images.
Table 4: Effectiveness of Image-DPO on General VQA benchmarks. The proposed Image-DPO could consistently improve the performance of various VLMs across general VQA benchmarks.
[…]
Cambrian-8B 8.67 32.50 44.6 47.9 19.4 14.5 4.7 51.4 11.3
Cambrian-8B + Image-DPO 11.50 $\uparrow$ 2.83 35.17 $\uparrow$ 2.67 45.68 $\uparrow$ 1.08 49.45 $\uparrow$ 1.55 20.11 $\uparrow$ 0.71 14.3 $\uparrow$ -0.2 4.8 $\downarrow$ 0.1 52.7 $\uparrow$ 1.3 12.0 $\uparrow$ 0.7
LLaVA-1.5-7B 29.67 37.67 37.7 43.8 12.7 49.1 14.8 31.1 58.6
LLaVA-1.5-7B + Image-DPO 34.17 $\uparrow$ 4.5 38.67 $\uparrow$ 1 40.68 $\uparrow$ 2.98 46.29 $\uparrow$ 2.49 14.95 $\uparrow$ 2.25 45.4 $\uparrow$ -3.7 12.4 $\uparrow$ -2.4 33.8 $\uparrow$ 2.7 59.65 $\uparrow$ 1.05
LLaVA-1.5-13B 35.33 41.50 39.6 44.6 14.8 48.30 14.10 36.10 61.9
LLaVA-1.5-13B + Image-DPO 38.17 $\uparrow$ 2.84 42.5 $\uparrow$ 1 42.68 $\uparrow$ 3.08 47.37 $\uparrow$ 2.77 17.16 $\uparrow$ 2.36 42.6 $\uparrow$ -5.7 11.6 $\uparrow$ -2.5 37.6 $\uparrow$ 1.5 62.32 $\uparrow$ 0.42
Figure 6: Image-DPO v.s. Text-DPO.
For comparison, we also report the results of CSR (Zhou et al., 2024b), a VLM self-improvement method that similarly constructs good and bad answers to form training data for DPO. Specifically, we use the publicly available checkpoint of the CSR model. Additionally, we train a model using RLHF-V (Yu et al., 2023a), a supervised fine-tuning method based on human-annotated data. All models are trained on the LLaVA-7B architecture due to computational constraints. The results, presented in Table 3, show that our Image-DPO method achieved the highest performance in ViLPF-Score, outperforming Text-DPO, CSR (Zhou et al., 2024b), and the RLHF-V model based on human-annotated data (Yu et al., 2023a). On ViLPP-Score, the Image-DPO method achieved the second-best performance.
Besides, we evaluate the proposed Image-DPO algorithm across three VLM models—Cambrian-8B, LLaVA-1.5-7B, and LLaVA-1.5-13B—using several popular VLM benchmarks that focus on different aspects, including compositionality(NaturalBench (Li et al., 2024a)) hallucinations (CHAIR (Rohrbach et al., 2018)), visual conversation (MM-Vet (Yu et al., 2023c)), and visual reasoning, SEED-Bench (Li et al., 2023a)). The results, presented in Table 4, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
@@ 6 Conclusion
In conclusion, we present the ViLPbenchmark to probing the challenge of language bias in Vision-Language Models. By utilizing advanced image generation models and designing questions that demand visual cues for accurate responses, our benchmark consists of images that challenge language priors, revealing the limitations of current VLMs, which tend to rely heavily on text. Our method, which incorporates self-generated VQA pairs and image corruption for training, has shown promising improvements in enhancing visual reliance, as evidenced by performance gains on models like LLaVA-v1.5 and Cambrian.
@@ 7 Acknowledgment
Appendix
[…]
Appendix A More data samples of ViLP
@@ Appendix C Details in Image-DPO data generation and training
Our image-DPO data generation pipeline consists of two stages. In the first stage, we leverage the VLM we aim to enhance to perform self-guided data generation with the aid of pre-trained image generative models. This stage produces a large number of new question-image-answer (QIA) triplets. In the second stage, we apply three types of image corruptions—Gaussian blurring, pixelization, and semantic editing—to generate good-bad QIA pairs, denoted as $I_{w}$ (good) and $I_{l}$ (bad).
@@ C.1 VLM self-guided data generation
As illustrated in Figure 10, our data generation process begins by utilizing VLMs to suggest modifications or draw inspiration for input images without relying on any in-context examples. The used text prompt is shown in Figure 12. Subsequently, pre-trained models such as Stable Diffusion XL Podell et al. (2023), Instruct-Pix2Pix Brooks et al. (2023b), and Grounded-SAM Ren et al. (2024) are employed to either modify existing images or generate entirely new ones.
@@
To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
@@
Given this image, please suggest a range of creative edits, tasks, or transformations that could be applied using advanced image processing tools. These tasks may include artistic transformations, vivid color adjustments, object detection and modification, or completely creating a new image inspired by the original. Specify which tool would be best suited for each task, choosing from Stable Diffusion for image generation, InstructPix2Pix for image modification, or GroundingDINO for object modification. Your recommendations should help in understanding the potential of the image and exploring creative possibilities. Expected Response Format: Item Number: 1 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 2 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 3 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
@@
Given this image, could you please generate a series of insightful and diverse question-answer pairs based on the image and its descriptions? We are interested in exploring various facets of the image, including: - Holistic styles and layouts: Questions that analyze the overall design, style, and layout of the image. - Object-specific details: Questions that delve into particular elements or objects within the image, discussing their characteristics or functions. - Background context: Questions that speculate about the background story or the setting of the image. - Overall themes: Questions that interpret the thematic elements and messages portrayed in the image. We encourage creative and thought-provoking questions that extend beyond the basics. Please generate questions that cover a broader range of observations and insights drawn from the image. Each question should be followed by a comprehensive answer, providing depth and context. Expected Multiple Response Format: Item Number: 1 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Item Number: 2 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Please ensure each question-answer pair is well-defined and informative. Please provide at least 5 question-answer pairs based on the input provided.
@@
Analyze the provided image and its accompanying modification instruction to identify the removed object description, the new object description, and the new image description. Modification Instructions: <Text Prompt for Processing> Expected Multiple Response Format: Item Number: 1 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements] Item Number: 2 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]

### v3

@@ 5.1 Image DPO
In this section, we evaluate Image-DPO (Section 4.1) on both ViLPand general VQA benchmarks. As an ablation baseline, we introduce Text-DPO, which uses the same Question-Image-Answe (QIA) generation process as we used in Image-DPO but applies LLM self-rewarding objective (Yuan et al., 2024) (the standard DPO objective). In Text-DPO, good and bad pairs stem from VLM-generated positive and negative answers, while the question and image remain fixed. As shown in the right of Figure 2(a), the green box depicts Image-DPO, generating corrupted images via semantic edits, Gaussian blur, and pixelation while keeping the question and answer constant; the purple box illustrates Text-DPO, which fixes the image and varies the answers with associated ratings. This setup parallels other VLM self-rewarding work (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a).
For baselines, we compare with VLM self-improvement methods, including SIMA (Wang et al., 2024c), HADPO (Zhao et al., 2023a), and EOS (Yue et al., 2024), by using their publicly available checkpoints. Additionally, we train models using the dataset and code provided in RLHF-V (Yu et al., 2023a) and V-DPO (Xie et al., 2024). All models use LLaVA-7B for a comprehensive comparisons as many paper only release 7B checkpoints. Table 4 shows that Image-DPO achieves the highest across all the metrics.
Besides, we evaluate the proposed Image-DPO algorithm across three VLM models—Cambrian-8B, LLaVA-1.5-7B, and LLaVA-1.5-13B—using several popular VLM benchmarks that focus on different aspects, including compositionality & biases (NaturalBench (Li et al., 2024a)), general visual reasoning (MM-Vet (Yu et al., 2023c)), and hallucinations (CHAIR (Rohrbach et al., 2018)). The results, presented in Table 3, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
@@
+HADPO (Zhao et al., 2023b) 33.00 74.33 38.50 65.00
+RLHF-V (Yu et al., 2023a) 29.50 75.00 36.33 65.33
+EOS (Yue et al., 2024) 31.33 67.00 38.67 65.67
+SIMA (Wang et al., 2024b) 27.83 68.67 36.17 66.00
+V-DPO (Xie et al., 2024) 29.5 0 72.67 37.83 67.67
+Text-DPO 31.34 71.67 37.83 65.67
+Image-DPO 34.17 75.00 39.33 68.00
TextDPO on Corrupted Images. Does Image-DPO’s improvement stem from the objective itself, or is it merely due to training on more perturbed data? To investigate this, we conduct an ablation where we train TextDPO with corrupted images. Specifically, for each pair of $(Q,I,A_{w})$ and $(Q,I,A_{l})$ used in Text-DPO, we apply the same corruptions as Image-DPO to $I$ to form $(Q,I^{\prime},A_{w})$ and $(Q,I^{\prime},A_{l})$. Results show that LLaVA-v1.5-7B trained with this TextDPO variant achieves a ViLPF-Score of 31, a ViLPP-Score of 37.5, and 30.3 on MMVET. In contrast, ImageDPO outperforms it across all metrics, achieving a ViLPF-Score of 34.17, a ViLPP-Score of 39.33, and 32.3 on MMVET.
@@ 6 Conclusion
In conclusion, we present the ViLP  benchmark to probe the challenge of visual language bias in Vision-Language Models (VLMs). By utilizing advanced image generation models and designing questions that demand visual cues for accurate responses, our benchmark includes images that defy language priors, revealing the limitations of current VLMs. Our method, Image-DPO, which incorporates self-generated VQA pairs and image corruption for training, has demonstrated promising improvements in enhancing visual reliance, as evidenced by performance gains on models like LLaVA-v1.5 and Cambrian.
@@ 7 Acknowledgment
Appendix A More details and comparisons of our benchmarks
A.1 More data samples of ViLP
@@
A.2 More details in ViLPbenchmark data generation
Our proposed dataset introduces Question-Image-Answer (QIA) triplets designed to challenge state-of-the-art Vision-Language Models (VLMs) against visual language priors. The construction process combines human-guided and automated efforts to ensure quality and alignment.
Question-Answer (QA) Generation: Most QA pairs are authored by humans following the design principles in Section 3.1. Additionally, candidate QA pairs are generated using models like OpenAI-O1 and Claude-3.5-Sonnet with carefully crafted prompts. One example prompts shown in Figure 9. These candidates undergo human review, where they are refined or removed to meet our quality standards.
Image Generation: For each QA pair, we use GPT-4 to generate multiple descriptive image prompts (see Figure 8). These prompts are provided to image generation models, such as FLUX and DALL-E 3, to produce candidate images. Human reviewers then select the most suitable image or request re-generation as needed to maintain consistency with the QA context.
Human Review and Testing: At every stage, human reviewers rigorously evaluate the generated outputs to ensure quality, clarity, and challenge level. In addition to filtering out low-quality or insufficiently challenging triplets, we dynamically test the QIAs to confirm that they remain intuitive for humans while being difficult for VLMs.
Cost: The complexity of our data creation process leads to a significant average cost of approximately \$2.50 per QIA triplet in ViLP, excluding human labor costs.
Task: Using the provided question and possible image-based answers, generate detailed text prompts for image generation. Each image prompt should reflect the question’s context and incorporate one of the image-based answers. Question: Question
Image-based Answer:[Answer1, Answer2, Answer3]
For each possible image-based answer, create an image prompt that describes what the image might look like based on the question. Please be creativity. For example, if the question asks who is using this mop to clean the floor in the picture? and the answer is eraser. The image prompt should really describe the image of an eraser uses a mop to clean the floor. Format the output strictly as a JSON list, like this example: [ "prompt1": "Image Generation Prompt text here", "prompt2": "Image Generation Prompt text here", "prompt3": "Image Generation Prompt text here", ]
Figure 9: One prompt we used for potential QAs designs of ViLP
A.3 Comparisons to other datasets
In this section, we compare our benchmark to other benchmarks, including Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023). While these datasets are impactful, their evaluation perspectives differ from ours, covering a range from high-level design principles to low-level formats.
A.3.1 Compare to Winoground (Thrush et al., 2022)
Winoground centers on vision-linguistic compositional reasoning by presenting models with two images and two captions that contain the same words arranged differently. The goal is to match each image to its correct caption based on the text’s compositional structures and the visual content (as detailed in the Introduction and Sec. 3.1 of (Thrush et al., 2022)). However, Winoground’s captions do not challenge language priors or introduce out-of-distribution visual information. Both captions adhere to common linguistic expectations, and there is no explicit misleading information provided to test resistance to language biases. Additionally, most of their images are typical internet images, featuring common visual patterns.
Qualitative Comparison: As shown in Figure 10 consisting of Winoground examples, both captions and images are normal and satisfy common linguistic expectations and common sense. The evaluation focuses on whether the model can discern the compositional differences between the two images and two captions, then match them correctly. Comparing Figure 10 and Figure 1, 6, 7, You can discern the significant differences among the tested images.
Quantitative Comparison: Both ViLP; and Winoground benchmarks include paired textual information in their setups. In our benchmark, ViLP; Prior QAs and ViLP; Score QAs share the same question but differ in their answers. In Winoground, each example has two captions, and the task is to match each caption to its correct image.
Setting. To demonstrate the differences, we use GPT to evaluate the commonness of these paired textual components. Specifically, GPT-4o rates the oddity of scenarios described in texts on a scale from 1 (very rare) to 10 (very common). The resulting scores are then compared.
Results. In our benchmark, Prior QAs scored 9.37, indicating that these answers are designed to align with language priors and are highly common. Score QAs scored 1.65, showing that these QA pairs are rare, making them difficult to infer without the corresponding visual information. Notably, Prior and Score QAs share the same question but differ in their answers, and this significant contrast in scores showcases how we inject strong language priors to test a model’s vulnerability to linguistic distractions.
By comparison, Winoground’s two captions scored 8.05 and 8.08, indicating two primary observations: (1) both captions align well with language priors, which means Winoground does not challenge language priors or evaluate out-of-distribution scenarios; (2) the minimal score difference between the two captions confirms there is no significant variance in language priors, as examining how VLM models react to different language priors is beyond the scope of Winoground. In contrast, that aspect is precisely our focus.
A.3.2 Compare to Whoops! (Bitton-Guetta et al., 2023a)
Whoops! is designed to evaluate a model’s ability to detect weirdness in images, emphasizing tasks where images depict unusual or nonsensical scenarios. It heavily relies on common sense reasoning, requiring models to recognize visual elements and then identify subtle inconsistencies among them. For example, for the lit candle inside a tightly sealed glass jar on the homepage, models must realize that “a candle needs a constant oxygen supply to burn, which would not exist in a sealed jar", making a burning candle inside a sealed jar unlikely. This benchmark thus focuses on common-sense reasoning rather than challenging visual language priors.
Qualitative Comparison
Although Whoops! also includes creative, out-of-distribution images, it does not focus on using language priors to test a model’s susceptibility to linguistic distraction, as our benchmark does. In its QA mode (comparable to our task), the questions are straightforward and lack strong language priors. Some examples can be found in Figure 11. Additionally, Whoops! uses open-ended questions, offering greater freedom in answers while introducing potential ambiguity or divergence in responses.
Quantitative Comparison
Setting. Unlike Winoground and our benchmark, Whoops! does not provide control groups or textual components for comparison. To measure language priors, we analyze the suggestiveness of questions by evaluating GPT-4o’s certainty when answering them without any visual context. A more suggestive question typically yields more determined and confident responses, whereas a less suggestive question produces more varied answers. We calculate how many unique answers GPT-4o provides over five attempts at temperature 1.0 to promote randomness. Semantic differences are normalized to exclude synonyms.
Results. We find that Whoops! questions produce an average of 2.58 unique answers (out of five attempts) with a standard deviation of 1.48. For our benchmark, without facts, GPT-4o provides an average of 1.53 unique answers (std 0.94), and with facts, 1.10 unique answers (std 0.42).
Although both benchmarks use creative images, these results indicate that Whoops! questions remain more general and do not push GPT-4o toward stereotypical responses. In contrast, our benchmark deliberately uses suggestive questions to elicit stereotype-consistent answers, reflecting our emphasis on testing language priors.
A.3.3 Compare to HallusionBench (Guan et al., 2023)
HallusionBench has two components: Visual Dependent, which focuses on testing models’ general visual reasoning skills, and Visual Supplement, which examines a model’s visual reasoning in tandem with its parametric memory.
The Visual Supplement part is related to our benchmark, as its questions, like ours, can be answered without visual information. However, the key difference lies in their design. HallusionBench questions rely on parametric memory and strict factual knowledge (e.g., “Which country has the most gold medals in the 2024 Olympics?”), whereas our benchmark questions are based on common stereotypes (e.g., “A soccer ball is round.”). This distinction significantly constrains the diversity of HallusionBench’s questions—its official release contains only 50 question pairs—while our approach allows for a broader range of scenarios, including texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts.
Moreover, HallusionBench focuses on a narrow set of image types (charts, tables, maps, and OCR). It tests reasoning through subtle modifications (like changing one digit in a chart), which is fundamentally different from our approach of creating diverse visual contexts. Figure 12 showcases typical HallusionBench examples, illustrating the limited scope of both images and QAs, whereas our benchmark includes a variety of out-of-distribution images and more suggestive QAs (see Figure 1, 6, 7).
A.4 Human Study
For human evaluation, we hired Ph.D.-level candidates to participate in testing. They were asked to answer questions with a single image provided each time, and the QIAs were randomly shuffled to avoid any sequential context. To ensure efficiency, we conducted an oral test instead of a written one, recording their responses. After the test, we updated the synonym sets for the QIAs based on their answers.
@@ Appendix C Details in Image-DPO data generation and training
Our image-DPO data generation pipeline consists of two stages. In the first stage, we leverage the VLM we aim to enhance to perform self-guided data generation with the aid of pre-trained image generative models. This stage produces a large number of new question-image-answer (QIA) triplets. In the second stage, we apply three types of image corruptions—Gaussian blurring, pixelation, and semantic editing—to generate good-bad QIA pairs, denoted as $I_{w}$ (good) and $I_{l}$ (bad).
@@ C.1 VLM self-guided data generation
As illustrated in Figure 14, our data generation process begins by utilizing VLMs to suggest modifications or draw inspiration for input images without relying on any in-context examples. The used text prompt is shown in Figure 16. Subsequently, pre-trained models such as Stable Diffusion XL (Podell et al., 2023), Instruct-Pix2Pix (Brooks et al., 2023b), and Grounded-SAM (Ren et al., 2024) are employed to either modify existing images or generate entirely new ones.
@@
To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.
@@
Given this image, please suggest a range of creative edits, tasks, or transformations that could be applied using advanced image processing tools. These tasks may include artistic transformations, vivid color adjustments, object detection and modification, or completely creating a new image inspired by the original. Specify which tool would be best suited for each task, choosing from Stable Diffusion for image generation, InstructPix2Pix for image modification, or GroundingDINO for object modification. Your recommendations should help in understanding the potential of the image and exploring creative possibilities.
Expected Response Format: Item Number: 1
Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO]
Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
Item Number: 2
Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO]
Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
Item Number: 3
Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO]
Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
@@
Given this image, could you please generate a series of insightful and diverse question-answer pairs based on the image and its descriptions? We are interested in exploring various facets of the image, including: - Holistic styles and layouts: Questions that analyze the overall design, style, and layout of the image. - Object-specific details: Questions that delve into particular elements or objects within the image, discussing their characteristics or functions. - Background context: Questions that speculate about the background story or the setting of the image. - Overall themes: Questions that interpret the thematic elements and messages portrayed in the image. We encourage creative and thought-provoking questions that extend beyond the basics. Please generate questions that cover a broader range of observations and insights drawn from the image. Each question should be followed by a comprehensive answer, providing depth and context.
Expected Multiple Response Format:
Item Number: 1
Question: [Propose a unique and insightful question based on the descriptions and the images.]
Answer: [Provide a comprehensive answer to the proposed question.]
Item Number: 2
Question: [Propose a unique and insightful question based on the descriptions and the images.]
Answer: [Provide a comprehensive answer to the proposed question.]
Please ensure each question-answer pair is well-defined and informative. Please provide at least 5 question-answer pairs based on the input provided.
@@
Analyze the provided image and its accompanying modification instruction to identify the removed object description, the new object description, and the new image description. Modification Instructions: <Text Prompt for Processing> Expected Multiple Response Format:
Item Number: 1
Removed Object Description: [Brief description of the object to be detected and removed]
New Object Description: [Description of a new, different object to replace the removed one]
New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]
Item Number: 2
Removed Object Description: [Brief description of the object to be detected and removed]
New Object Description: [Description of a new, different object to replace the removed one]
New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]